2. In a solution prepared by adding excess Pbl2(S) [Kgp = 1.4 x 10-8] to water, [1 ] at equilibrium is A) 8.4 x 10-5 mol/L. B is this answer e?
Answer #1

At equilibrium:

PbI2 <----> Pb2+ + 2 I-

   s 2s

Ksp = [Pb2+][I-]^2

1.4*10^-8=(s)*(2s)^2

1.4*10^-8= 4(s)^3

s = 1.5*10^-3 M

Now use:

[I-] = 2s

= 2*1.5*10^-3 M

= 3.0*10^-3 M

Answer: E